-t^2+3t+2=0
when you put x=1.5 why is it -(1.5)^2+3*1.5+2=4.25 instead of (-1.5)^2+4.5+2=8.75

Answer by Alan3354(69443)

-t^2+3t+2=0
when you put x=1.5 why is it -(1.5)^2+3*1.5+2=4.25 instead of (-1.5)^2+4.5+2=8.75
===================
-t^2 = -1*t^2, not the same as (-t)^2
----------
Same as at^2 = a*t^2 not (at)^2